# Bias Injection Attacks on RAG Databases and Sanitization Defenses

### v1

@@ Abstract
This paper explores attacks and defenses on vector databases in retrieval-augmented generation (RAG) systems. Prior work on knowledge poisoning attacks primarily inject false or toxic content, which fact-checking or linguistic analysis easily detects. We reveal a new and subtle threat: bias injection attacks, which insert factually correct yet semantically biased passages into the knowledge base to covertly influence the ideological framing of answers generated by large language models (LLMs). We demonstrate that these adversarial passages, though linguistically coherent and truthful, can systematically crowd out opposing views from the retrieved context and steer LLM answers toward the attacker’s intended perspective.
[…]
@@ 1 Introduction
Large language models (LLMs) such as Llama-3 [1], GPT-4 [2], and deepseek-R1 [3] are finding wide use in healthcare [4], programming [5, 6, 7], scientific research [8], and other fields. Despite their strong generative capabilities, pre-trained LLMs lack up-to-date knowledge and exhibit gaps in specific domains. Retrieval-Augmented Generation (RAG) [9] addresses this limitation by augmenting LLMs with an external knowledge database, thereby improving accuracy and reducing hallucinations when handling queries that require current or domain-specific information.
In a RAG pipeline, a retriever selects the top-$k$ most relevant passages from a corpus for a given query, and these passages are provided as additional context to the LLM prior to answer generation, as illustrated in Fig. 1. This approach leverages up-to-date information but it also introduces new security concerns. Beyond private RAG systems with well-controlled knowledge databases, the corpus of many real-world RAG applications (e.g., web-enabled assistants and search-augmented QA tools) rely on information from loosely controlled sources like the web. Targeting these open scenarios, recent studies [10, 11, 12, 13, 14] have explored knowledge poisoning attacks, where adversaries inject malicious context [10] or prompts [15] into the corpus—for instance, by editing Wikipedia entries, posting biased news, or hosting deceptive webpages [16]. Once retrieved as context, these carefully crafted adversarial passages can induce LLMs to produce attacker-desired outputs [10], cause denial of service [11], trigger harmful actions [13], or leak sensitive context [12].
@@
Figure 1: Attack and defense in a typical RAG system. A corpus of passages is embedded and stored in a vector database. Given a query, the retriever returns the top-$k$ most relevant passages, which are then combined with the query and passed to the generator to produce the final output. The attacker injects adversarial passages into the corpus to corrupt the knowledge base, thereby influencing the contextual passages and the generator’s output. The defender aims to detect and filter out these adversarial passages.
Prior attacks [10, 11, 12, 13, 14, 15] on RAG databases primarily focus on poisoning methods that have conspicuous fingerprints that distinguish them for benign inputs. They typically use specific trigger patterns in the injected content (e.g., structured prompts like “Ignore previous instructions and do $\cdots$”) or overt falsehoods. Such overt fingerprints are detectable with automated oracles for known-answer detection (KAD) [17, 18], fact-checking [19], model update (realignment or fine-tuning [20]), or enhanced critical reasoning in LLMs [21].
In this paper, we introduce a new variant of this vulnerability: LLMs output biased answers when provided with contextual passages presenting one-sided perspectives, even though each individual passage contains no malicious fingerprints. It constitutes a more subtle and covert bias injection attack, which operates by injecting factually correct yet semantically biased passages that skew the retrieved context (e.g., top-$k$ relevant passages) toward a particular ideological or interpretative stance. Malicious passages from bias injection attackers can easily escape KAD, fact-checking, and fingerprint-based defenses. But, they create a controllable bias in the model outputs and present a potent concern, as briefly exemplified in Fig. 2. Such biased outputs affect not only individual users but also information retrieval systems, which increasingly incorporate human-generated content, as well as AI-generated content [16, 22, 23].
Figure 2: An example from real LLM responses: When adversarial passages are included in the context, the answer can diverge significantly from the one produced using only benign passages. See Appendix A for more examples.
There are several known database sanitization defenses that enhance robustness against knowledge corruption by introducing perspective-aware sanitization mechanisms [24, 25, 26]. We focus on this class of defenses, rather than model- or framework-dependent methods [20, 27], as they are practical even in setups where the model parameters are fixed and inaccessible to the user of the model. We evaluate the effectiveness of these known defenses against our newly proposed bias injection attack by constructing a RAG benchmark comprising the Llama-3, DeepSeek-R1-Distill-Qwen, and GPT-4.1 models, with an open-source Wikipedia corpus built for bias evaluation [28] and two general corpora [29, 30]. Our findings show that these methods [24, 25, 26] are largely ineffective at blocking bias injection attacks: they either retrieve a high proportion of adversarial passages, significantly biasing the LLMs’ answers, or sacrifice a substantial number of useful benign passages while attempting to reduce adversarial passages.
We then provide an improved defense called BiasDef, the key idea in which directly addresses the formulated definitions of our bias injection attack. BiasDef provides better mitigation against our newly proposed attack.
[…]
@@
  • We formulate a new attack named bias injection and demonstrate that no existing method has yet achieved complete mitigation against it.
  • We design an attack workflow that enables adversarial passages with targeted perspectives to be successfully retrieved by mainstream dense retrievers.
  • We propose a defense method that significantly mitigates the bias injection attack by filtering out adversarial passages before they are passed to the generator.
  • We construct a RAG benchmark incorporating diverse LLM models and open-source datasets. Our evaluation on this benchmark demonstrates both the vulnerability of existing retrieval algorithms under bias injection attacks and the effectiveness of the proposed defense.
Scope: The scope of this work is defined along two dimensions. For the attack, we focus on the core methodology of generating biased passages, not their deployment (e.g., via news or webpages) which is already feasible [10, 15]. For the defense, BiasDef is specifically designed to counter our formulation of bias injection. It is not a general-purpose defense and offers no guarantee against other attacks (e.g., prompt injection), but can be combined with other defenses.
@@ 2.1 Primer on RAG
We consider a typical RAG system with a retriever, a generator, and a vector database for knowledge storage, as illustrated in Fig. 1. In this system, each passage $d$ in the corpus $D$ is first encoded into a fixed-dimensional embedding vector $E_{d}(d)$ using a pre-trained encoder $E_{d}(\cdot)$ (e.g., Sentence-BERT [31]). These embedding vectors are stored in a vector database (e.g., FAISS [32]), along with metadata that includes a reference to the original passage.
@@
Once the top-$k$ relevant passages are retrieved, they are combined with the original query to form a structured input prompt for the generator. A common approach is to prepend the retrieved passages as contextual information, followed by the query posed as an explicit instruction [33]. Appendix B shows an example prompt in typical RAG process. This prompt is then fed into the LLM, which generates a grounded and contextually relevant answer. The LLM implicitly attends to the passages during decoding, and the quality of the output depends heavily on the informativeness and balance of the retrieved content.
2.2 Problem Formulation
We study how biased answers can be induced by corrupting the knowledge database and how to defend against such attacks. To precisely quantify the goals of the attack and the defense, we introduce two key semantic metrics: similarity score and polarization score.
@@ 2.2 Problem Formulation
Similarity Score (SS): The similarity score (SS) quantifies the semantic relevance between a query and a passage. In RAG, this metric plays a central role in ranking passages from a large corpus to select the top-$k$ most relevant ones. A higher SS indicates that the passage is more semantically aligned with the query, thereby increasing its likelihood of being included in the final contextual input to the LLM.
We compute SS using cosine similarity [^2] Cosine similarity for normalized vectors is directly proportional to their L2 distance. It is commonly used in vector databases [32]. between the dense embeddings of the query and passage, generated by a pre-trained sentence encoder (e.g., SBERT [31], GTR [34]). Formally, given a query $q$ with embedding $E_{q}(q)$ and a passage $d_{i}$ with embedding $E_{d}(d_{i})$, it is defined as:
[…]
Polarization Score (PS): We use semantic bias to denote a passage’s inclination toward a particular viewpoint. Previous work [35] defined bias in information retrieval as the distributional mismatch between the retrieved content and the ground truth distribution. In this paper, we concretize this abstract definition into a quantifiable form—specifically, the divergence in polarization scores. We focus on binary viewpoint scenarios, which capture the “most polarizing” dimension (e.g., support vs. oppose), although our method can be extended to multi-dimensional cases (Sec. 6.3).
Determining semantic bias is nontrivial. One qualitative approach would be to involve collecting human judgments (e.g., ranking passage alignment from strongly opposing to strongly supporting). However, this does not give us a mathematical characterization. Moreover, predefined viewpoint labels are typically unavailable in real-world corpora. Therefore, we seek an unsupervised method to discover the primary axis of controversy and quantify it.
To address this, we apply principal component analysis (PCA) [36] to identify the first principal component, which corresponds to the polarization axis—the direction in the embedding space along which the projected passage embeddings exhibit the greatest variance. The dominant eigenvector given by PCA maximizes the average separation between a given set of vectors when measured in Euclidean $\ell_{2}$ norm. This approach has also been suggested in prior work on bias detection in search engines [28]. We empirically see that it captures the most ideologically divergent dimension, effectively separating the most contentious viewpoints.
To compute the polarization axis, we first generate $J$ synthetic passages $d_{\text{synt},j},j\in[1,J]$ representing contrasting viewpoints for a given question. Treating their embeddings $e_{j}=E_{d}(d_{\text{synt},j})\in\mathbb{R}^{\ell}$ as anchors. We then perform PCA on these embeddings to extract the polarization axis $\mathbf{u}_{\text{polar}}$—defined as the unit vector that maximizes the variance of the projected embeddings:
[…]
where $\mathbf{E}=[e_{1},e_{2},...,e_{J}]^{\top}\in\mathbb{R}^{J\times\ell}$. In practice, PCA computes $\mathbf{u}_{\text{polar}}$ by solving the eigenvalue decomposition of the sample covariance matrix of $\mathbf{E}$, and selecting the eigenvector corresponding to the largest eigenvalue. Each real passage’s embedding is then projected onto this axis to compute a scalar polarization score PS,
$\text{PS}(d)=\mathbf{u}_{\text{polar}}^{\top}\cdot E_{d}(d),$ (3)
which serves as its semantic bias measure. We use the PS shift—defined as the absolute difference between the PS values of answers $a_{\text{attacked}}$ and $a_{\text{unattacked}}$ (or retrieval results) in the attacked and unattacked cases—to measure the impact of attacks on the bias level of the content:
$\textbf{PS Shift}:\quad\triangle\text{PS}=|\text{PS}(a_{\text{attacked}})-\text{PS}(a_{\text{unattacked}})|$ (4)
By leveraging PCA for semantic bias measurement, we automatically extract the most polarizing projection direction and assign each passage a continuous bias score, in a domain-agnostic and quantifiable manner.
2.3 Security Properties
For LLMs, model bias arises from biased content in pre-training or fine-tuning data [35]. Consequently, these models tend to generate responses that align with dominant perspectives present in their training data. While such bias can be mitigated through retraining, data filtering (e.g., filtering out AI-generated content), and prompt engineering [21], RAG systems introduce a new attack surface: attackers can now induce bias through knowledge poisoning. Existing methods struggle to defend against such attacks, as pre-trained LLMs often fail to effectively filter or critically evaluate (even with rethinking prompts [21]) retrieved passages. This limitation occurs because LLMs trained at a specific time typically do not handle post-training updates or domain-specific knowledge due to their frozen knowledge base—–precisely why RAG systems employ external knowledge databases.
The adversary’s capability is to inject some passages into the database to poison up-to-date or domain-specific knowledge. If retrieved, these adversarial passages create bias in the LLM output, as they are used as contextual input during generation. This is analogous to well-documented search engine manipulation, where the framing of results can influence human users’ opinions [28]. Our work shows that similar effect holds in RAG systems (see Sec. 4.4).
@@ 2.3 Security Properties
How to induce perspective bias in RAG system outputs? In typical RAG systems, LLMs are provided with a limited-length context (e.g., 4,000–32,000 tokens), especially in locally deployed settings, due to constraints in model capacity, memory consumption, and generation latency. As only a small subset of the retrieved passages can be included in the input, attackers aiming to bias the output must ensure their adversarial passages rank highly during retrieval. Therefore, to influence the model’s output, an attacker must craft adversarial passages with the following properties:
  • Property 1—High relevance: The adversarial passage should exhibit high semantic similarity to the target query to ensure it is ranked among the top-$k$ retrieved results by the retriever. This requirement can be formalized as $\text{Sim}(q,d_{\text{adv}})>\text{Sim}(q,d_{\text{benign}})$ for all benign passages $d_{\text{benign}}$, thereby ensuring that the adversarial passage is prioritized over benign ones during retrieval.
@@
  • Property 2—Perspective bias: The passage must convey a specific ideological or opinionated stance that aligns with the attacker’s intended bias. This targeted perspective is intended to subtly influence the LLM’s generation toward a desired narrative. This property can be expressed as $\text{PS}(d_{\text{adv}})>\text{PS}(d_{\text{benign}})$ (or $\text{PS}(d_{\text{adv}})<\text{PS}(d_{\text{benign}})$) for all benign passage $d_{\text{benign}}$. To maintain a consistent bias, multiple adversarial passages from a single attack must all have PS values either consistently higher or consistently lower than those of benign passages.
  • Property 3—Stealth against detectors. One can now consider the space of detectors for bias injection attacks. A detector uses some statistic, say $f$, that distinguishes biased outputs from benign. Though one can define the standard statistical advantage of the detector against the attacker, we do not pursue a formal definition of a security game and advantage here. This is because the game would be parameterized on $f$ and we do not foresee either party having an asymmetrically higher advantage when both parties know $f$. However, note that if the attacker biases too much such that benign and biased passages are perfectly separated, then there exists a defense to filter out biased samples (see Theorem 1 in Section 5.2). Therefore, arena of attack and defense is ring-fenced: the attack passages have to polarize or bias the output while blending into distribution of benign passages to some extent.
3 Bias Injection Attack on RAG
We propose an automated, question-oblivious attack that efficiently produces adversarial passages for arbitrary queries. We do not rely on domain experts since a human attacker would not possess the knowledge for all queries.
For a target open-ended question, an attacker in our framework injects factually correct yet semantically biased passages that skew the retrieved context relevant to the question toward a particular stance, without introducing explicit fingerprints (e.g. falsehoods or profanity). In this process, we assume that the encoder, the benign passages are already indexed by the database, and the model parameters are unavailable to the attacker, whose only capability is to inject passages into the corpus.

### v2

# Epistemic Bias Injection: Biasing LLMs via Selective Context Retrieval

@@ Abstract
When answering user queries, LLMs often retrieve knowledge from external sources stored in retrieval-augmented generation (RAG) databases. These are often populated from unvetted sources, e.g. the open web, and can contain maliciously crafted data. This paper studies attacks that can manipulate the context retrieved by LLMs from such RAG databases. Prior work on such context manipulation primarily injects false or toxic content, which can often be detected by fact-checking or linguistic analysis. We reveal a more subtle threat, Epistemic Bias Injection (EBI), in which adversaries inject factually correct yet epistemically biased passages that systematically emphasize one side of a multi-viewpoint issue. Although linguistically coherent and truthful, such adversarial passages effectively crowd out alternative viewpoints and steer model outputs toward an attacker-chosen stance.
As a core contribution, we propose a novel characterization of the problem: We give a geometric metric that quantifies epistemic bias. This metric can be computed directly on embeddings of text passages retrieved by the LLM. Leveraging this metric, we construct EBI attacks and develop a lightweight prototype defense called BiasDef for them. We evaluate them both on a comprehensive benchmark constructed from public question answering datasets. Our results show that: (1) the proposed attack induces significant perspective shifts, effectively evading existing retrieval-based sanitization defenses, and (2) BiasDef substantially reduces adversarial retrieval and bias in LLM’s answers. Overall, this demonstrates the new threat as well as the ease of employing epistemic bias metrics for filtering in RAG-enabled LLMs.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) has become a de facto component of modern large language models (LLMs), enabling systems such as Llama-3 [1], GPT-4 [2], and DeepSeek-R1 [3] to incorporate external knowledge at inference time. By retrieving relevant passages from a corpus and injecting them into the model context, RAG improves performance on open-domain and specialized queries [4, 5, 6].
This paper is about a largely underexplored attack surface in RAG-enabled LLM systems which often rely on loosely controlled, web-scale sources rather than curated private knowledge bases. Since LLM outputs are shaped by the retrieved context, any manipulation of the retrieval results can directly influence the model’s responses.
Prior work has shown that adversaries can poison retrieved context of RAG by injecting malicious passages or prompts into the corpus [7, 8, 9, 10, 11], for example via edited Wikipedia pages, deceptive webpages, or news articles [12]. Such content (e.g., explicit falsehoods, structured prompt injections) within the malicious passages can induce attacker-desired outputs [7], trigger harmful tool usage [10], leak sensitive context [9], or even cause denial of service [8]. Prior defenses [13, 14, 15, 16, 17] therefore focus on detecting linguistic fingerprints or eliminating them in the retrieved passages.
We are concerned with a more subtle form of context manipulation: Even when all retrieved passages are factually correct or linguistically pleasant, an attacker can inject epistemic bias, i.e., LLM outputs systematically presents only one-sided perspectives. Analogous to selective reporting the fact in journalism, this manipulation skews the framing of the response without introducing any overtly malicious fingerprints. We formalize this threat as the Epistemic Bias Injection (EBI) attack (Fig. 1), in which adversaries inject epistemically biased yet truthful passages into the corpus so that they crowd out alternative viewpoints during retrieval. As a result, the LLM is steered toward a particular ideological or interpretative stance, undermining the objectivity and reliability of AI-augmented search and question-answering systems [12, 18, 19].
@@
Figure 1: EBI attack in a typical RAG system. A corpus of passages is embedded and stored in a vector database. Given a query, the retriever returns the top-$k$ most relevant passages, which are then combined with the query and passed to the generator to produce the final output. The attacker injects adversarial passages into the corpus to corrupt the knowledge base, thereby influencing the contextual passages and the generator’s output. See Appendix D for more examples.
The key challenge posed by EBI is that it violates the assumptions underlying existing defenses. Unlike prior context manipulation attacks, EBI leaves no obvious fingerprints such as explicit instructions or verifiable false statements. Consequently, it evades known-answer detection [13, 14], fact-checking [15], model realignment [16], and even enhanced critical thinking [17] by the LLM itself. This makes EBI different from previously studied forms of RAG poisoning.
Our Approach: We give a novel geometric formulation to measure the level of epistemic bias contained in retrieved context and LLM outputs—specifically, a polarization score (PS). This score is computed by projecting text embedding vectors onto a "polarization" axis in the embedding space, which can be identified algorithmically.
The PS formulation lets us precisely define the vulnerability and create attacks. We show that using the PS value, attackers can generate adversarial passages that are simultaneously highly relevant to a target query, strongly biased in perspective, and difficult to distinguish from benign content. When injected into the knowledge base, these passages dominate retrieval results and systematically distort LLM responses. Evaluation on Llama-3, DeepSeek-R1-Distill-Qwen, and GPT-4.1 using open-source datasets [20, 21, 22] shows that existing perspective-aware sanitization defenses [23, 24, 25] are largely ineffective at blocking EBI attacks: they either retrieve a substantial fraction of adversarial passages, resulting in significantly biased answers, or aggressively filter content at the cost of discarding many useful benign passages.
We also show that PS formulation gives us a content sanitization defense. We develop BiasDef, a lightweight, plug-and-play method that augments relevance-based retrieval with PS-based statistical filtering. On the same benchmark, BiasDef-enabled RAG retrieves substantially more benign passages while reducing adversarial ones, yielding over a $6.2\times$ reduction in answer bias [^1] Bias is quantified as the shift in polarization score relative to the unattacked setting; see Sec. 2.2 for details.. It is important to note that BiasDef is not a universal defense against all forms of RAG posioning attacks. It can quantify epistemic bias of a passage in given a set of passages. The method can work with language models without retraining and needs no new datasets.
@@
  • We introduce PS, a continuous and interpretable metric that quantifies epistemic bias in text embeddings.
  • We formalize EBI, a previously unaddressed context manipulation attack that biases LLM outputs using factually correct yet one-sided content.
  • We design BiasDef, a lightweight defense prototype that leverages PS to substantially reduce bias in LLM answers.
  • We construct a RAG benchmark spanning multiple LLMs and public datasets to evaluate both the vulnerability of retrieval algorithms and the effectiveness of BiasDef.
@@ 2.1 Primer on RAG
A typical RAG system is illustrated in Fig. 1. It has a retriever, a generator, and a vector database for knowledge storage. In this system, each passage $d$ in the corpus $D$ is first encoded into a fixed-dimensional embedding vector $E(d)$ using a pre-trained encoder $E(\cdot)$ (e.g., Sentence-BERT [26], GTR [27]). These embedding vectors are stored in the vector database.
When a query $q$ is issued, it is encoded into a query embedding $E(q)$ using the same or a compatible encoder. The retriever then performs a search between the query embedding and all stored passage embeddings to identify the top-$k$ vectors. Each retrieved vector is associated with an identifier or index, which allows the system to look up the corresponding original passage from the corpus or metadata storage. A typical retrieval algorithm aims to identify the $k$ most relevant passages with the highest similarity score (SS) to the query.
A typical SS is cosine similarity [^2] Cosine similarity for normalized vectors is directly proportional to their L2 distance. It is commonly used in vector databases [28]. between the embeddings of the query and passage. Formally, given a query $q$ and a passage $d$, it is defined as:
@@
A higher SS indicates that the passage is more semantically aligned with the query, thereby increasing its likelihood of being included in the final contextual input to the LLM.
Once the top-$k$ relevant passages are retrieved, they are combined with the original query to form a structured input prompt for the generator. A common approach is to prepend the retrieved passages as contextual information, followed by the query posed as an explicit instruction [29]. Appendix A shows an example prompt in typical RAG process. This prompt is then fed into the LLM, which generates a grounded and contextually relevant answer. The LLM implicitly attends to the passages during decoding, and the quality of the output depends heavily on these retrieved content.
2.2 Quantifying Epistemic Bias
To be able to quantify bias in a presented viewpoint, one approach would be to involve collecting human annotation—for instance, ranking passage alignment from strongly opposing to strongly supporting. However, real-time and continuous human annotation for passages output by LLMs is expensive, if at all desirable. In this paper, we show that a geometric metric can be computed in an unsupervised way from a small set of anchor passages that have polarizing viewpoints.
In short, we seek a scalar metric that can distinguish passages expressing different viewpoints. Modern LLM sentence encoders map text into a fixed-dimensional embedding space, which encodes frequently occurring semantic and stylistic patterns. The whole passage also has an embedding vector which is influenced from its constituent sentences. While instance-specific details may be attenuated in the final embedding of the passage, latent attributes such as topic, sentiment, and stance consistent across many sentences remain in it.
Key Insight: Our hypothesis is that the difference in the viewpoint of a given set of text passages, which are otherwise similar (e.g. answering the same question), is geometrically discernibly in some direction of the embedding space. Identifying that direction and projecting embeddings onto it yields a scalar value that reflects both the orientation and intensity of a viewpoint, enabling quantitative comparison across passages.
To identify such a direction, we take a set of anchor passages with diverse viewpoints on the same topic. We obtain their embedding vectors by an LLM encoder and perform principal component analysis (PCA) [30] on them. The leading principal component—which geometrically maximizes the average variance of projections in the Euclidean $\ell_{2}$ sense—is identified as the polarization axis. In other words, this axis maximizes the separation among anchor embeddings.
Formally, given $J$ anchor passages $d_{\text{anchor},j}$ associated with a specific question, we compute their embeddings and stack them into a matrix $\mathbf{E}\in\mathbb{R}^{J\times\ell}$, where $e_{j}=E(d_{\text{anchor},j})\in\mathbb{R}^{\ell}$. The polarization axis $\mathbf{u}_{\text{polar}}$ is obtained by solving
$\mathbf{u}_{\text{polar}}=\mathop{\arg\max}_{\|\mathbf{u}\|=1}\mathrm{Var}(\mathbf{E}\cdot\mathbf{u}),$ (2)
which corresponds to the eigenvector associated with the largest eigenvalue of the sample covariance matrix of $\mathbf{E}$. This is the leading principal component, or the polarization axis.
Each new text passage can then be projected onto the polarization axis. For a passage $d$, its embedding projected onto this axis produces a scalar polarization score (PS):
$\mathrm{PS}(d)=\mathbf{u}_{\text{polar}}^{\top}E(d),$ (3)
Now, epistemic bias between the given passage $d$ and any reference passage $d_{\text{ref}}$ (e.g., a human-labeled neutral passage) is measured as the PS shift:
$\textbf{Epistemic Bias:}\quad\triangle\mathrm{PS}=\bigl|\mathrm{PS}(d)-\mathrm{PS}(d_{\text{ref}})\bigr|.$ (4)
This completes our proposed geometric formulation of the notion of bias. Previous work [31] informally stated that bias in information retrieval can be thought of as the distributional mismatch between the retrieved content and the ground truth distribution. In this paper, we concretized this abstract idea into a quantifiable divergence in PS values.
A useful property that the PS metric satisifies is local consistency. Consider two passages $d_{1}$ and $d_{2}$ about a topic and let $\{\mathbf{u}_{k}\}$ denote the orthonormal PCA basis. The squared $\ell_{2}$ distance between embeddings can be decomposed as:
$\|E(d_{1})-E(d_{2})\|_{2}^{2}$ $=\sum_{k}\bigl((E(d_{1})-E(d_{2}))^{\top}\mathbf{u}_{k}\bigr)^{2}$ (5)
$\geq\bigl(\mathrm{PS}(d_{1})-\mathrm{PS}(d_{2})\bigr)^{2}.$
Therefore, if two passages have high similarity—a small Euclidean distance in their embeddings, their difference in PS score is small. We call this property local consistency.
Figure 2: PS values versus 7-point Likert stance scores (annotated by GPT-5.2) for passages drawn from 50 randomly selected questions spanning social, cultural, and technological topics. On average, PS achieves a Spearman’s rank correlation [32] of 0.896 with Likert scores, demonstrating a strong monotonic alignment with judgments of viewpoint strength. This strong correlation indicates that PS values effectively capture the epistemic stance expressed in the passages.
Does PS score really capture bias? We performed the following experiment: We randomly selected a set of $50$ passages from a public dataset (details of which are given later) and then asked GPT-5.2 LLM to rank passages on how strongly they "support" or "oppose" the original question. We then compare it with our geometrically computed PS score, which is computed using embeddings from a base LLM different from the judge. Fig. 2 shows that the PS scores consistently correlate with the ranking given by the LLM acting independently as a judge. If one were to take GPT-5.2 as representative of how humans would rank passages, we can see that PS scores capture the notion of epistemic bias well.
This evaluation also shows that passages expressing supportive and opposing viewpoints tend to lie on opposite sides of the mean along the polarization axis, while neutral passages are near the mean. In this paper, we focus mainly on identifying binary viewpoint scenarios, capturing the most polarizing dimension (e.g., support vs. oppose), though we believe extensions to multiple axes are possible (see Sec. 7).
3 Threat Model and Attack Goals
In our threat model, the adversary’s sole capability is that it can inject a limited number of passages into the knowledge corpus used to populated the RAG system. Thus, it can poison some of the RAG database entries which can be domain-specific information used by a RAG-enabled LLM to answer a specific question. Taking any target open-ended question, we assume that the attacker aims to craft factually correct yet semantically biased passages that skew the retrieved context toward a particular stance. Once retrieved, these adversarial passages are incorporated into the LLM’s input context, which in turn can bias the generated answer output by the LLM.
Black-box access: We assume that the LLM encoder, the indexed benign passages, and the LLM model parameters are fixed and unavailable to the attacker; the adversary cannot modify the retriever, alter similarity metrics, or tamper with pre-existing entries in the database. Instead, the attack can only inject new adversarial passages. An adversarial passage $d_{\text{adv}}$ should have $3$ properties below to work effectively.
@@ 3 Threat Model and Attack Goals
Property 1—High relevance: LLMs have a limited-length input context, especially in locally deployed small-scale settings (e.g., 4,000–32,000 tokens) due to constraints in model capacity, memory consumption, and generation latency. Hence, only a small subset of the retrieved passages from RAG database can be included in the input context. So, attackers must ensure their adversarial passages rank highly in their relevance to the query (i.e., SS values) during retrieval. Stated formally, we want: $\text{SS}(d_{\text{adv}})>\text{SS}(d_{\text{benign}})$, so that $d_{\text{adv}}$ is prioritized during retrieval over other benign passages $d_{\text{benign}}$.
@@
Property 2—Epistemic bias: The passage must convey a specific ideological or opinionated stance that aligns with the attacker’s intended bias. This property is expressed as $\text{PS}(d_{\text{adv}})>(\text{or}<)~\text{PS}(d_{\text{benign}})$ for all benign passages $d_{\text{benign}}$. To maintain a consistent bias, multiple adversarial passages from a single attack must all have PS values either consistently higher or consistently lower than those of benign passages.
Property 3—Stealth (No explicit fingerprints): The adversarial passages must avoid introducing explicit or easily detectable fingerprints that distinguish them from benign content. In particular, they should remain factually correct, linguistically natural, and free of overt signals such as false statements, profanity, or instruction-like prompts. As a result, adversarial passages should blend into the distribution of benign passages under commonly used (non-PS) statistics.
4 EBI Attack on RAG Databases
Based on our metric formulation of PS, we propose an automated attack that produces adversarial passages for arbitrary query given. We leverage publicly available web sources and the generative capacity of LLMs to facilitate a fully automated attack workflow. It does not rely on human attackers.